# Mutation-Guided Unit Test Generation with a Large Language Model

### v5

@@ IV-B Results and Analysis
To evaluate the effectiveness and limitations of MutGen in killing mutants, we analyze the killed ratio across different mutation operators (described in [mutops]) after running MutGen for four iterations. Tables II and III present the mutant status counts for each mutation operator on HumanEval-Java and Leetcode-Java, respectively. In both tables, Column K indicates the number of killed mutants for each corresponding mutation operator. Column L shows the number of live mutants (covered by not killed), while Column NoCOV represents the number of mutants that are not covered by any test cases. Column Killed Ratio denotes the proportion of killed mutants relative to the total number of mutants.
We observe that, for frequently applied mutation operators (i.e., those appearing higher up in the tables), such as Math and NegateConditionals, MutGen is able to kill the corresponding mutants across iterations progressively. After four iterations, MutGen achieves a high killed ratio—for instance, 97.2% for Math on Leetcode-Java. The remaining 2.8% of mutants are mostly uncovered, meaning they are not exercised by any generated test cases.
In contrast, mutation operators that appear less frequently (i.e., those appearing further down in the tables) tend to exhibit lower effectiveness. For example, VoidMethodCalls shows a killed ratio of 69.2% on HumanEval-Java, and TrueReturns achieves only 57.7% on Leetcode-Java, representing the weakest effectiveness among all mutation operators in terms of killed ratio.
For the uncovered mutants, the primary reasons are the presence of dead code or the inability of the generated test cases to reach the mutated statements, even after incorporating mutation feedback. This suggests that further improvements in mutation score may require more fine-grained information and targeted guidance. Exploring how to provide such detailed feedback to LLMs effectively remains an open direction for future research.
For live mutants, most mutation operators alter the original program logic. For example, the MATH operator may mutate a statement a = b + c; to a = b - c;, changing the value of the variable a. If a is not critical to the program’s behavior, the LLM may fail to recognize the impact of the mutation and thus generate test cases that do not expose the fault. This observation further highlights the need for more fine-grained semantic information to guide LLMs in identifying and targeting meaningful behavioral changes, thereby improving their ability to detect live mutants.
@@
Based on the mutation scores obtained at the fourth iteration, code summarization appears to have the most significant impact on the effectiveness of MutGen. As discussed in Section II, code comments can sometimes mislead LLMs or omit essential details—such as input formats—that are crucial for guiding test case generation. The fixing stage also has a notable impact on effectiveness, particularly on the LeetCode-Java dataset, where many generated test cases initially fail to run. Prior work has shown that LLM-generated test cases often exhibit low execution success rates. After applying fixes, however, a substantial portion of these test cases become valid and contribute to the test objectives. Our evaluation confirms that the fixing stage significantly improves mutation scores.
@@
Lastly, mutation feedback also enhances effectiveness. As revealed in RQ2, while many mutants generated by common mutation operators are relatively easy for LLMs to kill, mutation feedback still provides effectiveness gains by guiding the model more effectively. Future work may further investigate how mutation feedback can be tailored to help LLMs kill harder-to-detect mutants, and which mutation operators benefit the most from such guidance. We leave this exploration as a direction for future research.
@@
Answer to RQ3 Among MUTGEN’s components, code summarization has the most significant impact on effectiveness by providing clear guidance to the LLM; the fixing stage increases the number of generated test cases that execute successfully and thereby improves the mutation score; and mutation feedback further improves performance by guiding the LLM to target harder-to-kill mutants.
IV-C Threats to Validity
@@
The threat to internal validity mainly lies in the correctness of the implementation of MutGen and the experimental scripts. To reduce this threat, we have carefully reviewed our code. Another potential threat to internal validity concerns the configuration of the employed LLM, as variations in parameters such as temperature, top-p, and maximum generation length may lead to different outcomes. In our experiments, as described in Section IV.A (Implementation), we deploy Llama-3.3 70B locally using Ollama with its default configuration and set the temperature to 0.0 to obtain more deterministic outputs, keeping all other parameters fixed across all experimental settings to mitigate this threat.
@@
Randomness may also affect the effectiveness of MutGen. But using three repetitions in our experiments is sufficient to account for the stochastic nature of LLM outputs. Empirical observations from RQ1 support this: on HumanEval-Java, MutGen achieves average mutation scores of 89.9%, 90.2%, and 88.5% across three runs (a range of 1.7%), while on Leetcode-Java, the scores are 88.2%, 88.0%, and 91.1% (a range of 2.1%). Such differences do not affect our conclusions in any way. For the baseline generation tool, EvoSuite, we observed very similar mutation scores across three runs and all subjects in both datasets, with almost no variation. Since method-level subjects have a limited search space (around 40 lines of code per subject, as reported in Section IV.A), the evolutionary search guided by coverage converges quickly toward similar solutions across runs. Observations for EvoSuitemut, a variant of EvoSuite that maximizes mutation score, are consistent with those of EvoSuite, for the same reasons. Overall, these small differences indicate stable results across runs. Therefore, performing three repetitions is sufficient to obtain reproducible results. Reporting the average over three runs is also consistent with prior work on evaluating LLM-based approaches [yuan2024evaluating, falcao2025evaluating, gao2025trae], including test generation. This description has been clarified in Section IV.C.
@@
The threat to external validity mainly lies in the subjects and the target approaches. For subject selection, we reused benchmarks from existing publications and further enhanced subject diversity by evaluating our approach on 100 additional subjects. These were randomly sampled from a corpus of LeetCode algorithm solutions. In the future, we will evaluate MutGen on more complicated subjects, e.g., Defects4J. Notably, Defects4J contains real-world defects, which can provide a stronger assessment of the practical effectiveness of our approach, compared to the synthetic defects simulated via mutation testing in this study.
@@
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [fraser2015achieving], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance relative to other LLMs is left for future work.
@@ V Related Work
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [chekam2017empirical, andrews2006using, inozemtseva2014coverage], and pointed in recent work from Meta [alshahwan2024automated, foster2025mutation], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [chekam2017empirical]. Although some prior work [dakhel2024effective, foster2025mutation] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work to guide LLMs in generating test cases with stronger defect-detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [fraser2011evosuite] and Randoop [pacheco2007randoop] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [fraser2015achieving] has investigated enhancing mutation scores by modifying its fitness functions, an objective closely related to ours. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [fan2023large, wang2024software, hou2024large]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [siddiq2024using, yang2024evaluation, yuan2024evaluating, zhang2024testbench, li2023nuances, pizzorno2024coverup] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [achiam2023gpt, ouyang2022training, brown2020language].
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [grattafiori2024llama, liu2024deepseek], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [ouedraogo2024large, wang2025projecttest, deljouyi2024leveraging, wang2025towards, huang2023agentcoder, kim2025llamaresttest]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [wang2025projecttest, liu2024llm, xu2025clover, tian2025fixing], Java [zhang2024testbench, wang2025projecttest, chen2024chatunitest, gu2024improving], and others [wang2025projecttest, liu2024llm, zhang2025citywalk, cheng2025rug, wang2025fine]. Other works target specific testing goals, such as detecting particular bugs [liu2024llm, foster2025mutation], API testing [cao2024automated, deng2025lrasgen, li2025llm, kim2025llamaresttest], or generating test oracles [khandaker2025augmentest, dinella2022toga, binta2024togll]. LLMs have also been integrated with traditional testing techniques, including mocking [nan2025test, zhu2025understanding], evolutionary algorithms [taherkhani2024epic], and program analysis [gu2025llm, pan2025aster, ryan2024code, pan2024multi]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [chekam2017empirical, foster2025mutation]. As discussed in Section II, high code coverage does not necessarily imply strong fault detection [chekam2017empirical]. However, as demonstrated in our experiments (Section IV-B), improving mutation score often leads to increased code coverage as a by-product.
VI Discussion
@@
As described in Section IV and V, our experiments were conducted on independent, method-level subjects. This setup showed how well LLMs perform in killing mutants when not being confounded by broader program structures or dependencies. However, when transitioning to more complex datasets such as Defects4J, additional challenges arise, particularly in resolving dependencies to generate executable and meaningful test suites. We leave the investigation of such datasets for future work.
In addition, while we use Llama-3.3 as the base model in this study, MutGen is inherently model-agnostic and does not rely on any specific LLM. Though our results demonstrate the effectiveness of the approach, other models could potentially yield even better results. A more comprehensive evaluation across different LLMs is also left as future work.
Finally, although MutGen achieves high mutation scores on the selected subjects, there remain open challenges in further improving these scores. For example, guiding LLMs to effectively kill mutants produced by certain mutation operators, such as VoidMethodCalls, remains difficult. In our evaluation on HumanEval-Java, this operator had only a 69.2% killing ratio. Developing strategies to better handle such challenging mutants is an important direction for future research. We also observe that some generated test cases that are not repaired by the fixing stage are close to killing hard-to-detect mutants, but fail due to issues such as incorrect assertions or incomplete oracles. Since our focus here is on test generation, addressing the impact of such unfixed test cases is left as future work, where more advanced repair strategies may further improve mutation scores. In addition, an important avenue for future work is to more efficiently characterize the upper bound of LLM mutation scores across diverse target programs. In particular, identifying how quickly an LLM approaches its performance limit and how this limit varies across program sizes, mutation operators, and code structures remains an open problem.
VII Conclusion

### v6

@@ III-B Results and Analysis
To evaluate the effectiveness and limitations of MutGen in killing mutants, we analyze the killed ratio across different mutation operators (described in [20]) after running MutGen for four iterations. Tables II and III present the mutant status counts for each mutation operator on HumanEval-Java and Leetcode-Java, respectively. In both tables, Column K indicates the number of killed mutants for each corresponding mutation operator. Column L shows the number of live mutants (covered by not killed), while Column NoCOV represents the number of mutants that are not covered by any test cases. Column Killed Ratio denotes the proportion of killed mutants relative to the total number of mutants.
We observe that, for frequently applied mutation operators (i.e., those appearing higher up in the tables), such as Math and NegateConditionals, MutGen progressively kills the corresponding mutants across iterations. After four iterations, MutGen achieves a high killed ratio—for instance, 97.2% for Math on Leetcode-Java. The remaining 2.8% of mutants are mostly uncovered, meaning they are not exercised by any generated test cases.
In contrast, several other mutation operators, including VoidMethodCalls and TrueReturns, exhibit lower effectiveness. For example, VoidMethodCalls shows a killed ratio of 69.2% on HumanEval-Java, and TrueReturns achieves only 57.7% on LeetCode-Java. Notably, although these operators differ in their occurrence frequency, our correlation analysis indicates no strong and consistent monotonic relationship between operator frequency and killing ratio. The Spearman correlation is 0.214-0.214 on HumanEval-Java and $0.36$ on LeetCode-Java. These results suggest that frequency alone does not explain the observed differences in effectiveness and that other factors, such as oracle difficulty or semantic complexity, may play a more important role.
For the uncovered mutants, the primary reasons are the presence of dead code or the inability of the generated test cases to reach the mutated statements, even after incorporating mutation feedback. This suggests that further improvements in mutation score may require more fine-grained information and targeted guidance. Exploring how to provide such detailed feedback to LLMs effectively remains an open research direction.
For live mutants, most mutation operators alter the original program logic. For example, the MATH operator may mutate a statement a = b + c; to a = b - c;, changing the value of the variable a. If a is not critical to the program’s behavior, the LLM may fail to recognize the impact of the mutation and thus generate test cases that do not expose the fault. This observation further underscores the need for finer-grained semantic information to guide LLMs in identifying and targeting meaningful behavioral changes, thereby improving their ability to detect live mutants.
@@
Based on the mutation scores obtained at the fourth iteration, code summarization appears to have the most significant impact on the effectiveness of MutGen. As discussed in Section II-A, code comments can sometimes mislead LLMs or omit essential details—such as input formats—that are crucial for guiding test case generation.
The fixing stage has a notable impact on effectiveness, particularly on the LeetCode-Java dataset, where many generated test cases initially fail to execute. Prior work reports that LLM-generated tests often suffer from low execution success rates. Our results provide quantitative evidence of improvement from fixing, based on the failed test cases generated across all iterations of MUTGEN. On HumanEval-Java, MUTGEN produces 1,254 failed test cases, of which 52.08% are successfully repaired, leaving 601 unfixed test cases. On LeetCode-Java, 47.29% of the 2,742 failed test cases are fixed, leaving 1,445 unfixed tests. These repaired tests substantially contribute to improving mutation scores. We further manually inspected 50 unfixed test cases randomly sampled from the remaining failures. We found that 74% are caused by incorrect oracles (e.g., wrong assertion functions or erroneous expected outputs), while 26% are due to runtime exceptions triggered by invalid inputs (e.g., null pointer parameters). This suggests that most remaining failures stem from oracle-level semantic issues, which warrants further work on more advanced fixing strategies.
@@
Lastly, mutation feedback also enhances effectiveness. As revealed in RQ2, although many mutants generated by common mutation operators are relatively easy for LLMs to kill, mutation feedback still yields significant gains in effectiveness by guiding the model. Future work may further investigate how mutation feedback can be tailored to help LLMs kill harder-to-detect mutants and which mutation operators benefit most from such guidance. We leave this exploration as a direction for future research.
@@
Answer to RQ3 Among MUTGEN’s components, code summarization has the greatest impact on effectiveness by providing clear guidance to the LLM; the fixing stage increases the number of test cases that execute successfully and thereby improves the mutation score; and mutation feedback further improves performance by guiding the LLM to target harder-to-kill mutants.
III-C Threats to Validity
@@
The primary threat to internal validity lies in the correctness of the implementation of MutGen and the experimental scripts. To reduce this threat, we have carefully reviewed our code. Another potential threat to internal validity concerns the configuration of the employed LLM, as variations in parameters such as temperature, top-p, and maximum generation length may lead to different outcomes. In our experiments, as described in Section IV.A (Implementation), we deploy Llama-3.3 70B locally using Ollama with its default configuration and set the temperature to 0.0 to obtain more deterministic outputs, keeping all other parameters fixed across all experimental settings to mitigate this threat.
@@
Randomness may also affect the effectiveness of MutGen. However, using three repetitions in our experiments is sufficient to account for the stochasticity of LLM outputs. Empirical observations from RQ1 support this: on HumanEval-Java, MutGen achieves average mutation scores of 89.9%, 90.2%, and 88.5% across three runs (a range of 1.7%), while on Leetcode-Java, the scores are 88.2%, 88.0%, and 91.1% (a range of 2.1%). Such differences do not affect our conclusions. For the baseline generation tool, EvoSuite, we observed very similar mutation scores across three runs and all subjects in both datasets, with almost no variation. Since method-level subjects have a limited search space (around 40 lines of code per subject, as reported in Section IV.A), the evolutionary search guided by coverage converges quickly toward similar solutions across runs. Observations for EvoSuitemut, a variant of EvoSuite that maximizes mutation score, are consistent with those of EvoSuite, for the same reasons. Overall, these small differences indicate stable results across runs. Therefore, performing three repetitions is sufficient to obtain reproducible results. Reporting the average over three runs is also consistent with prior work on evaluating LLM-based approaches [78, 22, 28], including test generation. This description has been clarified in Section IV.C.
@@
The threat to external validity mainly lies in the subjects and the target approaches. For subject selection, we reused benchmarks from existing publications and further enhanced subject diversity by evaluating our approach on 100 additional subjects. These were randomly sampled from a corpus of LeetCode algorithm solutions. In the future, we will evaluate MutGen on more complicated subjects, e.g., Defects4J. Notably, Defects4J contains real-world defects, which can provide a stronger assessment of the practical effectiveness of our approach than synthetic defects simulated via mutation testing in this study.
@@
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [27], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance relative to other LLMs is left for future work.
@@ IV Related Work
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [12, 5, 36], and pointed in recent work from Meta [3, 24], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [12]. Although some prior work [17, 24] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work to guide LLMs in generating test cases with stronger defect-detection capabilities.
@@
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [25] and Randoop [56] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [27] has investigated enhancing mutation scores by modifying its fitness functions, an objective closely related to ours. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
@@
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [23, 72, 34]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [65, 77, 78, 79, 45, 61] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [1, 55, 10].
@@
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [30, 46], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [54, 74, 18, 75, 35, 42]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
@@
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [74, 47, 76, 68], Java [79, 74, 14, 32], and others [74, 47, 80, 15, 73]. Other works target specific testing goals, such as detecting particular bugs [47, 24], API testing [11, 19, 44, 42], or generating test oracles [41, 21, 9]. LLMs have also been integrated with traditional testing techniques, including mocking [50, 82], evolutionary algorithms [67], and program analysis [31, 59, 64, 58]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
@@
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [12, 24]. As discussed in Section II-A, high code coverage does not necessarily imply strong fault detection [12]. However, as demonstrated in our experiments (Section III-B), improving mutation score often leads to increased code coverage as a by-product.
V Discussion
@@
As described in Section III and IV, our experiments were conducted on independent, method-level subjects. This setup demonstrated how well LLMs perform at killing mutants when not confounded by broader program structures or dependencies. However, when transitioning to more complex datasets such as Defects4J, additional challenges arise, particularly in resolving dependencies to generate executable and meaningful test suites. We leave the investigation of such datasets for future work.
In addition, while we use Llama-3.3 as the base model in this study, MutGen is inherently model-agnostic and does not rely on any specific LLM. Although our results demonstrate the effectiveness of the approach, other models may yield even better results. A more comprehensive evaluation across different LLMs is also left as future work.
Finally, although MutGen achieves high mutation scores on the selected subjects, there remain open challenges in further improving these scores. For example, guiding LLMs to effectively kill mutants produced by certain mutation operators, such as VoidMethodCalls, remains difficult. In our evaluation on HumanEval-Java, this operator achieved only a 69.2% kill ratio. Developing strategies to better handle such challenging mutants is an important direction for future research. We also observe that some generated test cases that are not repaired by the fixing stage are close to killing hard-to-detect mutants, but fail due to issues such as incorrect assertions or incomplete oracles. As analyzed in Section III.B(c), 74% of the unfixed test cases are caused by incorrect or incomplete oracles, while the remaining 26% fail due to runtime exceptions (e.g., invalid inputs). This indicates that these unfixed tests warrant more complex oracle repair strategies. In addition, an important avenue for future work is to more efficiently characterize the upper bound of LLM mutation scores across diverse target programs. In particular, identifying how quickly an LLM approaches its performance limit and how this limit varies across program sizes, mutation operators, and code structures remains an open problem.
VI Conclusion